### v1

# Identifying Drift, Diffusion, and Causal Structure from Temporal Snapshots

Abstract
Stochastic differential equations (SDEs) are a fundamental tool for modelling dynamic processes, including gene regulatory networks (GRNs), contaminant transport, financial markets, and image generation. However, learning the underlying SDE from observational data is a challenging task, especially when individual trajectories are not observable. Motivated by burgeoning research in single-cell datasets, we present the first comprehensive approach for jointly estimating the drift and diffusion of an SDE from its temporal marginals. Assuming linear drift and additive diffusion, we prove that these parameters are identifiable from marginals if and only if the initial distribution is not invariant to a class of generalized rotations, a condition that is satisfied by most distributions. We further prove that the causal graph of any SDE with additive diffusion can be recovered from the SDE parameters. To complement this theory, we adapt entropy-regularized optimal transport to handle anisotropic diffusion, and introduce APPEX (Alternating Projection Parameter Estimation from $X_{0}$), an iterative algorithm designed to estimate the drift, diffusion, and causal graph of an additive noise SDE, solely from temporal marginals. We show that each of these steps are asymptotically optimal with respect to the Kullback–Leibler divergence, and demonstrate APPEX’s effectiveness on simulated data from linear additive noise SDEs.
1 Introduction
This work presents the first comprehensive approach for jointly estimating the drift and diffusion of a stochastic differential equation (SDE) from observed temporal marginals. While parameter estimation has been studied extensively using trajectory data, either given one long trajectory [12, 56, 9, 81, 53, 39, 57, 17], or multiple short trajectories [48, 87, 43, 67, 55], it is often impossible to observe individual trajectories when tracking large populations. For example, in single-cell RNA sequencing (scRNA-seq) datasets, the destructive nature of sequencing technologies prevents specific cells from being tracked over time [82, 80, 28, 78, 47]. Similarly, hydrogeochemical sensors cannot track distinct particles when studying contaminant flow, providing only data on plume migration [54, 37, 75, 24, 1, 13, 49].Observational data may therefore be limited to samples from temporal marginals of the process at various measurement times. We refer to this observational setting as marginals-only.
However, many distinct stochastic processes are known to be indistinguishable under the marginals-only setting [91, 18]. Due to these non-identifiability concerns, previous works simplify the problem significantly, such that at least one of diffusion or drift are already known [47, 74, 84, 20, 93, 27]. Moreover, it is common to impose model restrictions, such as irrotational drift and isotropic diffusion [47, 20, 93, 74, 84, 91]. In this work, we prove in Theorem 4.3 that non-identifiability under the time-homogeneous linear additive noise SDE model is resolved if the initial distribution $X_{0}$ is not auto-rotationally invariant (see Definition 4.2). We further introduce a practical method for estimating the drift and diffusion of a general additive noise process, without prior knowledge or additional assumptions, and extensively demonstrate its effectiveness on simulated data.
1.1 Background and related work
SDEs have long been used to model natural processes [62, 21, 42, 78], where the drift and diffusion have meaningful physical interpretations. In single-cell biology, the drift and diffusion model cell differentiation, where a stem cell changes into another cell type, such as a blood cell or bone cell. This process is often visualized with Waddington’s epigenetic landscape, where cells figuratively traverse a developmental terrain, guided by deterministic forces (drift) and stochastic fluctuations (diffusion) [85, 86, 78]. In particular, the drift can be related to the set of genes that directly regulate the expression of a gene of interest [93, 5, 4, 94, 83], and therefore provides insights into genetic diseases and gene therapies. Meanwhile, the diffusion informs the extent to which cell fates are determined by initial conditions [27]. In hydrological systems, drift is linked to important and unknown properties of the subsurface, such as hydraulic conductivity through average flow velocity [37, 8], while the diffusion usually describes material heterogeneity or turbulence [8], and therefore informs the applicability of popular models, such as Darcy’s law [54, 62, 44]. Together, these parameters determine pollutant fates in hydrological systems, a key concern in safeguarding drinking water sources [60, 26, 40, 16, 64, 15]. In some models, drift and diffusion are also directly dependent on one another, as in the chemical Langevin equation [30] and the advection-dispersion equation for solute transport [37, 44, 8]. We remark that even when the drift and diffusion are independent, these parameters remain inextricably linked, as they jointly explain the dynamics of the underlying process.
Inference from temporal marginals. The marginals-only observational setting dates back to at least Aristotle, who inferred the processes and mediators (e.g., heat) behind embryogenesis by dissecting bird eggs at various stages of development [85, 59, 3]. Most contemporary research in this area now focuses on single-cell biology. Many works focus on trajectory inference [78, 47, 92, 20], while others infer the causal graph, e.g. GRN [5, 10, 83, 70], or perform parameter estimation [19, 73]. Recent research jointly estimates a subset of these quantities by leveraging their intrinsic relationships. For instance, [84, 74] are primarily interested in trajectory inference, and iteratively estimate drift and trajectories. Meanwhile, [93] iteratively estimates drift and trajectories, while additionally applying permanent interventions on the drift dynamics for network inference.
Due to the difficulty of the problem, most approaches assume that the underlying diffusion is known [47, 74, 84, 20, 93]. Alternatively, [27] estimates diffusion, but assumes that an accurate estimate of drift is known. However, knowing drift or diffusion a priori is unrealistic in practice. In these settings, an accurate guess of one or the other parameter set is crucial because misspecified diffusion typically leads to poor drift estimation and vice-versa [8]. It is also standard to assume that the drift is a conservative field $\nabla\psi$ [91, 47, 84, 20, 74], and that the diffusion is given by isotropic Brownian motion $\sigma\mathrm{d}W_{t}$ [91, 47, 84, 20, 93, 74]. However, these constraints impose unrealistic conditions on the data. A conservative drift field cannot model negative feedback loops or repressilator dynamics [91], and isotropic diffusion cannot model different noise scales across variables or latent confounders [70, 51]. Finally, it is common to leverage additional perturbational data from interventions [70, 93], or impose regularization [93] to learn the system more completely. In contrast, given an initial distribution, which is not auto-rotationally invariant, we show that we can jointly estimate the drift, diffusion, and causal graph of any linear additive noise SDE, solely from observational temporal marginals.
Identifiability. Previous assumptions on prior knowledge and the underlying model are largely attributed to sources of non-identifiability introduced by the marginals-only observational setting [91, 18, 47, 27]. Establishing sufficient conditions for identifiability are therefore required to ensure that inference is feasible. [87] showed that identifiability of a time-homogeneous linear additive noise SDE from ground truth trajectories with fixed $X_{0}\in\mathbb{R}^{d}$ is equivalent to a non-degenerate rank condition based jointly on $X_{0}$, and the SDE parameters, given by the linear drift matrix $A$ and the observational diffusion $GG^{\top}$. In contrast, our work studies identifiability of linear additive noise SDEs in the more general setting of observing temporal marginals. As detailed in Section 3, the set of non-identifiable systems within our observational setting is strictly larger. Furthermore, we demonstrate that identifiability for linear additive noise SDEs can be guaranteed with a wide class of initial distributions. To the best of our knowledge, this has not been shown even in the case of trajectory-based observations.
By connecting SDE parameter identification to recent work on dynamic structural causal models [11], we show that full causal identification is possible from the linear additive noise SDE model, if the initial distribution $X_{0}$ is not auto-rotationally invariant. This implies non-Gaussianity of the initial marginal, which is reminiscent of linear non-Gaussian acyclic models, where the full causal graph is identifiable [76].
1.2 Our contributions
In this work, we show that it is possible to infer the drift, diffusion, and underlying causal graph, solely from appropriately initialized temporal marginals. We illustrate our theoretical contributions in Figure 1(a) and our algorithmic contributions in Figure 1(b). We further outline our contributions below said figures.
Temporal marginals observed froma non-auto-rotationally invariant $X_{0}$round truth SDE parametersCausal graph $\mathcal{G}$XTheorem 4.3Lemma 4.5
(a) Theoretical contributions: identifying drift, diffusion, and causal structure from temporal marginals
EstimateSDE parametersEstimatelaw on paths XAEOTMLE
(b) Algorithmic contributions: APPEX algorithm
Figure 1: Outline of our theoretical (a) and algorithmic (b) contributions. In (b), our algorithm Alternating Projection Parameter Estimation from $X_{0}$ alternates between Anisotropic Entropy-regularized Optimal Transport (AEOT) and Maximum Likelihood Estimation (MLE).
  1. 1.
    We provide a theoretical foundation for system identification from temporal marginals in Section 4.
    1. (a)
      In Theorem 4.3, we provide a full characterization of the identifiability conditions for the drift and diffusion parameters from temporal marginals, under the time-homogeneous linear additive noise model. Identifiability is guaranteed if and only if the initial distribution $X_{0}$ is not auto-rotationally invariant (see Definition 4.2).
    2. (b)
      In Lemma 4.5, we connect parameter estimation to causal structure learning via dynamic structural causal models [11]. We show that the causal graph of any additive noise model can be recovered from the SDE parameters.
  2. 2.
    In Section 5, we introduce our algorithm: Alternating Projection Parameter Estimation from $X_{0}$ (APPEX). It is the first method designed to estimate drift and diffusion from temporal marginals of additive noise SDEs, without prior knowledge.
    1. (a)
      We show that with each iteration, APPEX’s estimates approach the true solution. APPEX alternates between a trajectory inference step, which is a Schrödinger bridge problem, and a parameter estimation step, which is solved via maximum likelihood estimation. We show that both subprocedures are optimal with respect to the Kullback-Leibeler (KL) divergence.
    2. (b)
      To solve the associated Schrödinger bridge problem for trajectory inference, given arbitrary additive noise, we adapt entropic optimal transport for parameter dependent transport costs, including anisotropic diffusion.
    3. (c)
      We test APPEX’s efficacy across a wide range of experiments in Section 6. APPEX can identify the drift and diffusion parameters from the previously non-identifiable examples introduced in Section 3, if given appropriate non-auto-rotationally invariant initial distributions. We also demonstrate APPEX’s effectiveness over a large-scale experiment on higher dimensional SDEs with randomly generated parameter sets. Our results demonstrate that APPEX identifies arbitrary linear additive noise SDEs with significantly higher accuracy than the widely-used Waddington-OT (WOT) method [78]. Our final experiments demonstrate that APPEX can be used to identify the system’s causal graph, even in the presence of latent confounders.
2 Mathematical setup
Let $b(X_{t})$ be a Lipschitz function $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$, $G\in\mathbb{R}^{d\times m}$, and denote the standard $m$-dimensional Brownian motion as $W_{t}\in\mathbb{R}^{m}$. $X_{t}\in\mathbb{R}^{d}$ is the solution to the corresponding time-homogeneous additive noise SDE if
$\mathrm{d}X_{t}=b(X_{t})\>\mathrm{d}t+G\>\mathrm{d}W_{t},\ X_{0}\sim p_{0}.$ (1)
The condition that $b(X_{t})$ is a Lipschitz drift function ensures existence and uniqueness of the SDE solution [25]. The additive noise is modeled by the diffusion matrix $G\in\mathbb{R}^{d\times m}$. In this paper, we also refer to $H=GG^{\top}\in\mathbb{R}^{d\times d}$ as the (observational) diffusion, since it is only possible to observe $H$ [87, 65].
Time-homogeneous linear additive noise SDEs are among the most studied SDEs, and have been applied to numerous fields of interest [87]. These SDEs generalize the popular Ornstein-Uhlenbeck process, such that stationarity is not assumed. A time-homogeneous linear additive noise SDE obeys the form
$\mathrm{d}X_{t}=AX_{t}\>\mathrm{d}t+G\>\mathrm{d}W_{t},\ \ X_{0}\sim p_{0},\ \ A\in\mathbb{R}^{d\times d}$ (2)
The primary objective of this paper is to estimate the drift and diffusion parameters of the SDE from temporal marginals. Given an SDE, a temporal marginal is the distribution of the process at time $t$, which we denote by $p_{t}\sim X_{t}$. In particular, given $N$ observed measurement times $t_{0}=0,\ldots t_{N-1}$, we would like to estimate the drift and diffusion of the SDE given the set of measured marginals $\{p_{i}\}_{i=0}^{N-1}$. For a linear additive noise SDE, this corresponds to identifying $A$ and $H=GG^{\top}$ from $\{p_{i}\}_{i=0}^{N-1}$.
The Fokker-Planck equation (3) describes the continuous evolution of the process’ temporal marginals $p_{t}$ from the initial distribution $p_{0}$. A straightforward computation shows that the marginals of a linear additive noise SDE evolve according to the following Fokker-Planck equation:
$\frac{\partial}{\partial t}p(x,t)=-\nabla\cdot(Ax)p(x,t)+\frac{1}{2}\nabla\cdot(H\nabla p(x,t)),\ \ p(x,0)=p_{0}(x).$ (3)
Although (3) assumes differentiability, we note that it can also be defined in the weak distributional sense for general probability measures $p_{t}$ (see Appendix Section 8.2).
In the finite sample setting, we observe $M_{i}$ samples for the $i$th marginal. The empirical marginal distribution is given by $\hat{p}_{i}\sim\mathrm{Unif}\{x_{t_{i}}^{(j)}:j\in 1,\ldots M_{i}\}$, where $x_{t_{i}}^{(j)}$ is the $j$th observation of the process at time $t_{i}$. If we take the number of observations $M_{i}$ per time to be infinite, then our observed marginals $\hat{p}_{i}$ would equal the true marginal distributions $p_{i}$ of the SDE.
Notation Meaning
$b(X_{t})$ Drift function of a general SDE
$A$ Drift matrix of a linear SDE
$G$ Diffusion matrix of an additive noise SDE
$H=GG^{T}$ Observational diffusion matrix
$p_{t}\sim X_{t}$ Temporal marginal distribution of the SDE at time $t$
$p(x,t)$ Probability density of the process at time $t$, assuming existence
$X_{i}=X_{t_{i}}$ Distribution of SDE at $i$th observation time
$\Delta X_{i}$ $X_{i+1}-X_{i}$
$p_{i}\sim X_{t_{i}}$ Temporal marginal distribution of the SDE at measured time $t_{i}$
$N$ Number of observed temporal marginals
$\hat{p}_{i}$ Empirical temporal marginal of the SDE at measured time $t_{i}$
$x_{t_{i}}^{(j)}$ $j$th sample from $i$th empirical marginal
$M_{i}$ Number of samples for the $i$th empirical temporal marginal $\hat{p}_{i}$
$\Sigma$ Covariance of an initial distribution $X_{0}$
$\epsilon$ Scalar parameter for regularizing entropy-regularized optimal transport (EOT)
$\Sigma(\theta)$ (Possibly) anisotropic matrix parameter for regularizing
anisotropic entropy regularized optimal transport (AEOT)
[…]
3 Examples of non-identifiability
The problem of non-identifiability given $p_{0}$ appears if there exists an alternative drift-diffusion pair, $(\tilde{A},\tilde{H})\neq(A,H)$, which shares the same time marginals, following initial distribution $X_{0}\sim p_{0}$. Equivalently, the processes share the same Fokker-Planck equation (3) across all observed times. We first summarize a few classical examples of non-identifiability from the literature [47, 74, 87, 33, 91]. In all examples, the processes $X_{t}$ and $Y_{t}$ share the same time marginals.
Example 3.1 (Starting at stationary distribution [47, 74]).
This example is motivated by the fact that multiple SDEs may share the same stationary distribution. Hence, if $p_{0}$ is initialized at the stationary distribution, then each of these SDEs will be non-identifiable from their marginals. We provide a simple example.
$\mathrm{d}X_{t}$ $=-X_{t}\>\mathrm{d}t+\mathrm{d}W_{t},$ $X_{0}\sim\mathcal{N}(0,\frac{1}{2})$ (4)
$\mathrm{d}Y_{t}$ $=-10Y_{t}\>\mathrm{d}t+\sqrt{10}\mathrm{d}W_{t},$ $Y_{0}\sim\mathcal{N}(0,\frac{1}{2})$ (5)
Here, both SDEs have the same stationary distribution $\mathcal{N}(0,\frac{1}{2})$ despite having significantly different individual trajectories. Indeed, the stationary distribution of a $1$-dimensional 00-mean Ornstein-Uhlenbeck (OU) process with drift $-\lambda X_{t}$ and diffusion $\sigma$ is $\mathcal{N}(0,\frac{\sigma^{2}}{2d\lambda})$, which depends only on the drift:diffusivity ratio $\lambda/\sigma^{2}$ [47]. For multivariate OU processes with drift $A$ and observational diffusion $H$, the stationary distribution $\mathcal{N}(0,\Sigma)$ depends only on the relationship $\Sigma A+A\Sigma=-H$ [52]. Intuitively, the set of OU processes, which share the same stationary distribution $\mathcal{N}(0,\Sigma)$ are equivalent up to rescaling time.
Example 3.2 (Rotation around process mean [74, 33, 91]).
In this example, the first SDE governing $X_{t}$ is $2$-dimensional Brownian motion and the second SDE governing $Y_{t}$ adds a divergence-free rotational vector field $(x,y)\to(y,-x)$ about $(0,0)$, which is undetectable if the initial distribution $p_{0}$ is rotationally invariant.
$\mathrm{d}X_{t}$ $=\mathrm{d}W_{t},$ $X_{0}\sim\mathcal{N}(0,Id)$ (6)
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}Y_{t}\>\mathrm{d}t+\mathrm{d}W_{t},$ $Y_{0}\sim\mathcal{N}(0,Id)$ (7)
Hence, if $p_{0}$ is an isotropic distribution with mean $(0,0)$, then $X_{t}$ and $Y_{t}$ are non-identifiable from one another [74]. This can also be shown directly with the Fokker-Planck equation (3), since $\nabla\cdot(Ax)p(x,t)=\nabla p(x,t)\cdot Ax+p(x,t)\nabla\cdot A(x)=\nabla p(x,t)\cdot Ax=0$, if $p$ is parallel to the rotational vector field $Ax=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}x$.
Example 3.3 (Degenerate rank [87]).
In this example motivated by [87], trajectories (and therefore temporal marginals) of distinct SDEs are non-identifiable if the process is degenerate, limiting observations to a strict subspace of $\mathbb{R}^{d}$.
$\mathrm{d}X_{t}$ $=\begin{bmatrix}1&2\\ 1&0\end{bmatrix}X_{t}\>\mathrm{d}t+\begin{bmatrix}1&2\\ -1&-2\end{bmatrix}\mathrm{d}W_{t},$ $X_{0}=\begin{bmatrix}1\\ -1\end{bmatrix}$ (8)
$\mathrm{d}Y_{t}$ $=\begin{bmatrix}1/3&4/3\\ 2/3&-1/3\end{bmatrix}Y_{t}\>\mathrm{d}t+\begin{bmatrix}1&2\\ -1&-2\end{bmatrix}\mathrm{d}W_{t},$ $Y_{0}=\begin{bmatrix}1\\ -1\end{bmatrix}$ (9)
The drift matrices $\begin{bmatrix}1&2\\ 1&0\end{bmatrix},\begin{bmatrix}1/3&4/3\\ 2/3&-1/3\end{bmatrix}$ each have eigenvector $X_{0}=\begin{bmatrix}1\\ -1\end{bmatrix}$ with eigenvalue $-1$. Moreover, the diffusion is rank-degenerate with column space $span(\begin{bmatrix}1\\ -1\par\end{bmatrix})$. Thus, both SDEs will have identical behaviour along $span(\begin{bmatrix}1\\ -1\par\end{bmatrix})$ and are only differentiated by their behaviour elsewhere. Given $X_{0}=\begin{bmatrix}1\\ -1\end{bmatrix}$, the processes will stay within $span(\begin{bmatrix}1\\ -1\par\end{bmatrix})$ and are non-identifiable from one another. We note that this non-identifiability holds even when we observe trajectories [87], whereas the first two examples are identifiable from trajectories but not identifiable from marginals.
In order to contend with these sources of non-identifiability, the standard assumptions are that the drift is an irrotational vector field and that the diffusion is both known and isotropic [91, 47, 20, 74]. The primary goal of Section 4.1 is to show when the SDE parameters are identifiable without such assumptions. We discuss the restrictive implications of previous assumptions in Section 4.2.
4 System identification from temporal marginals
We provide a full characterization of the conditions for identifiability of linear additive noise SDEs from observed time marginals in Section 4.1. In particular, identifiability is determined by generalized rotational properties of the initial distribution $X_{0}$, for which we provide preliminary definitions below. We then show in Section 4.2 that one can recover the underlying causal graph from the identified parameters of a general additive noise SDE.
Definition 4.1.
We define a $\Sigma$-generalized rotation in $\mathbb{R}^{d}$ as the matrix exponential $e^{At}$, such that $A\in\mathbb{R}^{d\times d}$ is skew-symmetric with respect to $\Sigma\succeq 0$, i.e. $A\Sigma+\Sigma A^{\top}=0$. In particular, if the rotation is applied to a r.v. $X$ with covariance $\Sigma$, then the covariance will be preserved, i.e. $\Sigma=Cov(X)=Cov(e^{At}X)$ (Lemma 8.5 in Appendix).
$\Sigma=\gamma\mathrm{I}_{d}$ defines classical rotations, such that spheres, e.g. $S^{d-1}=\{x\in\mathbb{R}^{d}:\gamma x^{\top}x=1\}$, are rotationally invariant. For anisotropic choices of $\Sigma$, ellipsoids, e.g. $E_{\Sigma}=\{x\in\mathbb{R}^{d}:x^{\top}\Sigma^{-1}x=1\}$, are rotationally invariant. If $\Sigma\succ 0$, then $\Sigma$-generalized rotations can be interpreted as classical rotations within the $\Sigma$-weighted inner product space $\mathbb{R}^{d}_{\Sigma}$, since they preserve the norm $\|x\|_{\Sigma}^{2}=x^{\top}\Sigma^{-1}x$.
Definition 4.2.
Let $X$ be a $d$-dimensional r.v. with covariance $\Sigma$. We say that $X$ is auto-rotationally invariant, if there exists a nontrivial $\Sigma$-generalized rotation $e^{At}$ with $A\neq 0$ such $e^{At}X\sim X$ $\forall t\geq 0$. Conversely, $X$ is not auto-rotationally invariant, if $e^{At}X\sim X$ $\forall t\geq 0$ admits only the solution $A=0$.
Auto-rotational invariance can be interpreted as a random variable being rotationally invariant under the geometry induced by its own covariance $\Sigma$. For example, Gaussians $\mathcal{N}(0,\Sigma)$ and uniform distributions over ellipsoids are auto-rotationally invariant random variables. Additionally, all rank degenerate r.v.’s are auto-rotationally invariant, since $e^{At}X\sim X$ would be satisfied by any matrix $A$, such that $X$ is in its nullspace with probability $1$. However, given full-dimensional support, auto-rotational invariance requires strict ellipsoidal symmetry on the r.v.’s density (Lemma 8.6 in Appendix), which means that almost all non-degenerate r.v’s are not auto-rotationally invariant. We provide more examples and theory about rotations in Appendix 8.3, and also refer the reader to [63] for additional theory about elliptical rotations.
4.1 SDE parameter identifiability
We are now ready for our main identifiability theorem, which characterizes identifiability of linear additive noise SDEs.
Theorem 4.3.
Let $X_{t}$ evolve according to a $d$-dimensional time-homogeneous linear additive noise SDE (2), with initial condition $X_{0}\sim p_{0}$, such that all the moments of $p_{0}$ are finite. Then, the drift $A$ and the diffusion $GG^{\top}$ of the SDE can be uniquely identified from the time marginals $p_{t}$ if and only if $X_{0}\sim p_{0}$ is not auto-rotationally invariant.
Proof.
We present the proof in the case where $p_{t}$ has smooth density $p(x,t)$. The analogous argument for the case where $p_{t}$ is any probability measure is included in Appendix 8.3.
We first show that if $X_{0}$ is auto-rotationally invariant, then there exist multiple processes with distinct drift-diffusion parameters $(A,H=GG^{\top})$, which would share the same time marginals $p_{t}$, when initialized at $X_{0}\sim p_{0}$. Let $X_{0}$ be an auto-rotationally invariant r.v. with $cov(X_{0})=\Sigma$. Then, there exists a matrix $A\neq 0$ such that $e^{As}X_{0}\sim X_{0}$ $\forall s\geq 0$. To define the non-identifiable SDEs, the idea is to generalize the non-identifiable isotropic rotation example (3.2). In particular, we will define the non-identifiable SDEs such that only one of the SDEs has nonzero drift $A$, and both SDEs have diffusion proportional to $GG^{\top}=\Sigma$. We will prove that for all $\gamma\in\mathbb{R}$, the SDEs
$\mathrm{d}X_{t}$ $=\gamma G\mathrm{d}W_{t}$ (10)
$\mathrm{d}X_{t}$ $=AX_{t}\>\mathrm{d}t+\gamma G\mathrm{d}W_{t}$ (11)
will have the same time marginals $p(x,t)$. The Fokker-Planck equations for the two SDEs are given by
$\frac{\partial}{\partial t}p(x,t)$ $=\frac{\gamma^{2}}{2}\nabla\cdot(\Sigma\nabla p(x,t))$ $p(x,0)=p_{0}(x),$ (12)
$\frac{\partial}{\partial t}p(x,t)$ $=-\nabla\cdot(Ax)p(x,t)+\frac{\gamma^{2}}{2}\nabla\cdot(\Sigma\nabla p(x,t))$ $p(x,0)=p_{0}(x).$ (13)
To show that the processes exhibit the same time marginals, we show that (12) and (13) are equivalent by showing that the first divergence term in (13) is identically zero. We begin by showing equivalence at $t=0$. We need to show that for all $x\in\mathbb{R}^{d}$,
$\nabla\cdot(Ax)p(x,0)=\nabla p(x,0)\cdot Ax+p(x,0)\nabla\cdot(Ax)=0.$ (14)
By the auto-rotational invariance of $X_{0}$, $p(x,0)$ satisfies: $p(e^{As}x,0)=p(x,0)\quad\forall s$. Taking the derivative with respect to $s$ at $s=0$ and using chain rule, we have:
$0=\frac{\mathrm{d}}{\mathrm{d}s}p(x,0)\bigg{|}_{s=0}=\frac{\mathrm{d}}{\mathrm{d}s}p(e^{As}x,0)\bigg{|}_{s=0}=\nabla p(x,0)\cdot\frac{\mathrm{d}}{\mathrm{d}s}e^{As}x\bigg{|}_{s=0}=\nabla p(x,0)\cdot Ax,$
so the first term on the right side of (14) vanishes. We now show $\nabla\cdot(Ax)=\text{tr}(A)=0$ to prove (14). Lemma 8.5 implies that $A\Sigma+\Sigma A^{\top}=0$ and Lemma 8.9 shows that without loss of generality, we may pick $A$ so that $Tr(A)=0$. This proves (14) for $t=0$. The same argument applies for $t>0$ upon noticing that for both systems, $X_{t}$ is also auto-rotationally invariant with respect to the same map $e^{As}X_{t}\sim X_{t}$. This is shown in the Appendix under Lemma 8.12. This demonstrates that if $X_{0}$ is auto-rotationally invariant, then we do not have identifiability from marginals over the class of linear additive noise SDEs.
We now prove that identifiability is guaranteed if $X_{0}$ is not auto-rotationally invariant. Suppose that two linear additive noise SDEs with parameters $(A,H=GG^{\top})$ and $(\tilde{A},\tilde{H}=\tilde{G}\tilde{G}^{\top})$ have the same time marginals $p(x,t)$ given the same initial distribution $X_{0}\sim p_{0}$, which is not auto-rotationally invariant. We will prove that $(A,H)=(\tilde{A},\tilde{H})$ must hold. Since both SDEs share the same marginals, their Fokker-Planck equations are equivalent:
$\frac{\partial}{\partial_{t}}p(x,t)$ $=L(p(x,t)):=-\nabla\cdot(Ax)p(x,t)+\frac{1}{2}\nabla\cdot(H\nabla p(x,t))$ $p(x,0)=p_{0}(x)$ (15)
$=L_{1}(p(x,t)):=-\nabla\cdot(\tilde{A}x)p(x,t)+\frac{1}{2}\nabla\cdot(\tilde{H}\nabla p(x,t))$ $p(x,0)=p_{0}(x)$ (16)
By linearity, we may subtract both equations and obtain
00 0 $=(L-L_{1})(p(x,t))=-\nabla\cdot(\bar{A}x)p(x,t)+\frac{1}{2}\nabla\cdot(\bar{H}\nabla p(x,t))$ $p(x,0)=p_{0}(x),$ (17)
where $\bar{A}=A-\tilde{A}$ and $\bar{H}=H-\tilde{H}$. Hence, $p_{0}$ is a stationary solution to (17), such that $p_{t}=p_{0}$ for all $t\geq 0$ Now, (17) is precisely the Fokker-Planck equation for the residual SDE, which is the linear additive noise SDE given by
$\mathrm{d}X_{t}=\bar{A}X_{t}\>\mathrm{d}t+\bar{G}\mathrm{d}W_{t},$ (18)
such that $\bar{G}\bar{G}^{\top}=\bar{H}$. $X_{t}\sim X_{0}$ $\forall t\geq 0$. Then, we note that if $\bar{H}\neq 0$, (18) can only admit a Gaussian stationary distribution. In particular, stationarity would require that (18) is an Ornstein-Uhlenbeck process [23], whose stationary distribution is $\mathcal{N}(0,\Sigma)$, where $\Sigma\bar{A}+\bar{A}\Sigma=-\bar{H}$ [52]. Thus, if $X_{0}$ is not auto-rotationally invariant, then $X_{0}$ is non-Gaussian, and we must conclude that $\bar{H}=H-\tilde{H}=0$. Furthermore, this implies that the residual SDE is deterministic, such that
$\frac{\mathrm{d}X_{t}}{\mathrm{d}t}$ $=\bar{A}X_{t}\implies X_{t}=e^{\bar{A}t}X_{0}$ (19)
$X_{t}$ $=e^{\bar{A}t}X_{0}\sim X_{0}\hskip 14.22636pt\forall t\geq 0.$ (20)
However, note that $e^{\bar{A}t}X_{0}\sim X_{0}$ admits only the trivial solution $\bar{A}=0$ since $X_{0}$ is not auto-rotationally invariant. This proves that $A=\tilde{A}$ and $H=\tilde{H}$ as desired. ∎
Theorem 4.3 tells us that identifiability from marginals is guaranteed if the initial distribution $X_{0}$ is not auto-rotationally invariant. Conversely, if $X_{0}$ is auto-rotationally invariant, then it is possible to construct multiple SDEs with the same ensuing marginals. However, we note that in practice, observed marginals may not be compatible with our constructed example (11). An open question is if non-identifiability persists given any set of observed marginals, which start from an auto-rotationally invariant distribution $X_{0}$. As discussed previously, most distributions are not auto-rotationally invariant, and are hence conducive to identifiability. Figure 2 demonstrates how non-auto-rotationally invariant initial distributions $X_{0}$ resolve the non-identifiability of temporal marginals from the examples in Section 4. We discuss this in Appendix 8.4.
(a) This corresponds to Example 3.1 with $X_{0}\sim\delta_{1}$. The SDE with drift $A=H=10$ (right) has converged to the stationary distribution (red) at time $t=0.2$, while the SDE with $A=H=1$ (left) has not yet sufficiently drifted or diffused.
(b) This corresponds to Example 3.2, with $X_{0}\sim\text{Unif}\{(2,0),(2,0.1)\}$. The first SDE is irrotational and maintains mean $(2,0.05)$. The second SDE (right) rotates about the origin. At time $t=1.5$, the mean has shifted to approximately $(0,-2)$.
(c) This corresponds to Example 3.3, with $X_{0}\sim\text{Unif}\{(1,0)^{\top},(0,1)^{\top}\}$. The first SDE (left) has eigenvalue $2$ for $(2,1)$, while the second (right) has eigenvalue $1$, leading to more extreme intercepts in the left plot at $t=0.8$.
Figure 2: Marginals are plotted at various times for the SDE pairs from Section 4 following various non-auto-rotationally invariant initial distributions.
4.2 Causal graph identification
Given that the drift and diffusion are identified, we can gain important insights into the system’s causal structure. It has already been shown that knowing drift and diffusion provides the system’s post-intervention distributions [34, 87]. We will prove that under basic conditions, we can recover the causal graph $\mathcal{G}$ of an additive noise SDE from the drift $b(X_{t})$ and observational diffusion $H=GG^{\top}$.
We first review how a causal graph can be defined from an SDE. Consider a $d$-dimensional additive noise SDE (1). By isolating the $j$th variable and integrating, we obtain:
$X_{t}^{(j)}$ $=X_{0}^{(j)}+\int_{0}^{t}b_{j}(X^{\alpha(j)}_{s})\,\mathrm{d}s+\sum_{k=1}^{m}G_{j,k}W_{t}^{(k)},$ (21)
where $W_{t}$ is an $m$-dimensional Brownian motion, $\alpha(j)\subset[d]$ is the set of variables that are used as input in $b_{j}(X_{t})$. Then, (21) is an example of a dynamic structural causal model [11]. In particular, we can define a causal graph $\mathcal{G}$ over vertices $V=[d]$ and edges $E=\{i\to j:j\in[d],i\in\alpha(j)\}$ [11][Definition 2]. Intuitively, given $j\in V$, we include the edge $i\to j$ if $X^{(i)}$ influences the evolution of variable $X^{(j)}$ through the drift. We note that in non-additive noise models, edges $i\to j$ can also be formed due to the diffusion. In this paper, we we assume that latent confounders are modeled with shared driving noise.
Definition 4.4.
Let $X$ evolve according to a $d$-dimensional additive noise SDE: $\mathrm{d}X_{t}=b(X_{t})\>\mathrm{d}t+G\>\mathrm{d}W_{t}$. We define a multi-directed edge $\tilde{e}=(i_{1},\ldots i_{p})\in\mathcal{G}$ if variables $X^{(i_{1})},\ldots,X^{(i_{p})}$ depend on a common noise source $W_{t}^{(k)}$, $k\in[m]$. To explicitly represent the latent confounder, one may also use the alternative graphical representation of adding the vertex $U_{i_{1},\ldots i_{p}}$, with directed edges pointing to each of $i_{1},\ldots,i_{p}$ (as done in [35]).
This definition is motivated by [46], where multidirected edges represent sets of variables, which share a hidden latent confounder. We also note that it is common to model latent confounders with correlated diffusion in the dynamic setting (see [45, 51], and [11][Example 4]).
Lemma 4.5.
Let $X_{t}$ evolve according to a $d$-dimensional additive noise SDE: $\mathrm{d}X_{t}=b(X_{t})\>\mathrm{d}t+G\mathrm{d}W_{t}$, and let $\mathcal{G}=(V=[d],E,\tilde{E})$ be its causal graph, over the variables $X^{(1)},\cdots,X^{(d)}$, with directed edge set $E$ and multidirected edge set $\tilde{E}$.
  1. a.
    There exists a directed edge $e=i\to j$ in $\mathcal{G}$ if and only if $b_{j}(X_{t})$ depends on $X_{t}^{(i)}$.
  2. b.
    There exists a multidirected edge $\tilde{e}$ containing $(i,j)$ in $\mathcal{G}$, if and only if $H_{i,j}=(GG^{\top})_{i,j}\neq 0$.
Thus, $\mathcal{G}$ is fully identified by $b(X_{t})$ and $H=GG^{\top}$ if no more than two observed variables share a common noise source.
Proof.
The first claim follows immediately from the definition of the dynamic structural causal model (21).
To prove the second claim, it is easy to see from (21) that $X^{(i_{1})},\ldots,X^{(i_{p})}$ depends on a common noise source if and only if $G_{i_{1},k},\ldots G_{i_{p},k}\neq 0$ for some column $k\in[m]$ in the diffusion $G$. Indeed, the existence of such a column implies that $X^{(i_{1})},\ldots,X^{(i_{p})}$ each depend on $W_{t}^{(k)}$, and the other direction is given by the fact that $W_{t}$ is a multivariate Gaussian process, with independent components. If two variables share a noise source, this condition is equivalent to $H_{i,j}=G_{i}\cdot G_{j}\neq 0$. We therefore conclude that there is a multidirected edge containing $(i,j)$ in $\mathcal{G}$ iff $H_{i,j}\neq 0$.
Although the observational diffusion $H=GG^{\top}$ informs the presence of latent confounders between pairs of variables, we note that it is impossible for $H$ alone to determine general multidirected edges $(i_{1},\cdots,i_{p})$ over $p>2$ components, which indicate an unobserved confounder of $X_{t}^{(i_{1})},\cdots,X_{t}^{(i_{p})}$. Indeed, given only $H=GG^{\top}$, multiple causal interpretations may be possible, because structurally different matrices $G,\tilde{G}$ can obey $H=GG^{\top}=\tilde{G}\tilde{G}^{\top}$. We provide an example in Example 8.1 of the Appendix, where we cannot distinguish between a single multidirected edge $(1,2,3)$, and three pairs of bidirected edges $(1,2),(1,3),(2,3)$. Furthermore, we note that shared driving noise is not a comprehensive characterization of latent confounders in the dynamic setting, since confounding effects cannot always be absorbed into the diffusion. For an in depth overview of general dynamic structural causal models and their relation to defining causal graphs, we refer the reader to Appendix 8.1 and references [34, 51, 11].
Despite these limitations, Lemma 4.5 highlights the power of drift-diffusion identification towards recovering the causal graph. Indeed, the proposition accommodates any stochastic process with additive noise, and is thus equipped to handle complex deterministic relationships between variables, including feedback loops. This setting is particularly well suited for inferring GRNs, which are commonly modeled via additive noise SDEs [91, 90, 70, 93, 47, 20].
The causal framework also highlights the impracticality of common model assumptions within the scRNA-seq literature, namely isotropic diffusion $H=GG^{\top}=\sigma^{2}\mathrm{I}_{d}$ and irrotational drift $b(X_{t})=\nabla\psi$. We have seen in Lemma 4.5 that latent confounders from shared noise can only be modeled with anisotropic diffusion. Moreover, while we expect GRNs to contain feedback loops, i.e. cycles in $\mathcal{G}$, imposing irrotational drift prevents most cycles from being considered. We show this for the linear additive noise case below.
Lemma 4.6.
Let $X_{t}$ evolve according to a linear additive noise SDE (2). Then, the drift $AX_{t}$ is irrotational if and only if $A$ is symmetric, i.e. $A=A^{\top}$.
Proof.
First, we note that for any symmetric matrix $A$, the vector field $Ax$ can be expressed as the gradient of the scalar potential given by the quadratic form $\phi(x)=x^{\top}Ax/2$ [69][(96)-(97)]. Hence, $Ax$ is irrotational if $A=A^{\top}$.
For the opposite direction, we recall the fact that any gradient field from $\mathbb{R}^{n}\to\mathbb{R}^{n}$ has a symmetric Jacobian [89, 14]. Since the Jacobian of a linear vector field $Ax:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is $A$, we conclude that all irrotational linear vector fields from $\mathbb{R}^{n}\to\mathbb{R}^{n}$ are given by symmetric matrices. ∎
Lemma 4.6 shows that imposing irrotationality on a linear model is equivalent to imposing a symmetric drift, which determines a symmetric causal graph. However, as illustrated in [91][Fig. 6], symmetric GRNs cannot capture important relationships, such as negative feedback loops or repressilator dynamics. Indeed, consider the $2$-cycle $i\leftrightarrow j$. By Lemma 4.5, $i\to j$ if $A_{j,i}\neq 0$ and $j\to i$ if $A_{i,j}\neq 0$. However, if the model is irrotational, then by Proposition 4.6, $A_{i,j}=A_{j,i}$, which can only model a positive feedback loop.
To conclude this section, it is important to note that combining Theorem 4.3 and Lemma 4.5 shows that we can identify the full causal structure (e.g., GRN) if and only if $X_{0}$ is non-auto-rotationally invariant.
5 Our parameter estimation method
In this section, we introduce our method, Alternating Projection Parameter Estimation from $X_{0}$ (APPEX), which estimates a process’ drift, diffusion, and causal graph from observed temporal marginals. Figure 3 illustrates the general strategy of the APPEX algorithm. We focus on the case where the noise $GdW_{t}$ is additive and the drift $b(X_{t})=AX_{t}:A\in\mathbb{R}^{d\times d}$, is linear (2), since, under these conditions, Theorem 4.3 guarantees identifiability given that $X_{0}$ is not auto-rotationally invariant. However, our algorithm is applicable to any parametric drift family. In this section, we give an intuitive explanation at the population level, using the process’ true temporal marginals $p_{i}\sim X_{t_{i}}$. In Section 6, we perform experiments using empirical marginals $\hat{p}_{i}$, with finite samples per marginal.
Given observed temporal marginals $p_{0},...,p_{N-1}$ and an initial guess for drift and diffusion $(A^{(0)},H^{(0)})$ , the idea is to use an alternating optimization algorithm to obtain increasingly better estimates for the drift and diffusion matrices. Formally, we consider the spaces
$\mathcal{D}$ =$=\{q:q_{0}=p_{0},...,q_{N-1}=p_{N-1}\}\text{ and }$ (22)
$\mathcal{A}$ =$=\{p:\exists(A,H)\in\mathbb{R}^{d\times d}\times S^{d}_{\succeq 0}\text{ s.t. }p(x,t|y,s)\sim\mathcal{N}(e^{A(t-s)}y,H(t-s))\},$ (23)
where $\mathcal{D}$ is the set of laws on paths, which coincide with the $N$ temporal marginals at their respective times, and $\mathcal{A}$ is the set of laws given by linear additive noise SDEs. To find the underlying linear additive noise SDE in the intersection $\mathcal{A}\cap\mathcal{D}$, our method will alternate between information projections, $\operatorname*{arg\,min}_{q\in\mathcal{D}}D_{KL}(q\|p)$ and moment projections, $\operatorname*{arg\,min}_{p\in\mathcal{A}}D_{KL}(q\|p)$. Intuitively, the information projection step corresponds to trajectory inference, whose goal is to find the law on paths $q\in\mathcal{D}$, which best aligns with the reference measure $p$. The moment projection then corresponds to maximum likelihood parameter estimation, since it amounts to determining the law on paths $p\in\mathcal{A}$ over linear additive noise SDEs, which best aligns with the law of the estimated trajectories $q$. Indeed, in the infinite-data limit, MLE is equivalent to minimizing relative entropy, i.e. KL divergence [2, 88]. At the population level, an iteration entails the updates:
$q^{(k)}$ $=\operatorname*{arg\,min}\limits_{q\in\mathcal{D}}D_{KL}(q\|p_{A^{(k-1)},H^{(k-1)}})$ (24)
$A^{(k)}$ $=\operatorname*{arg\,min}\limits_{A\in\mathbb{R}^{d\times d}}D_{KL}(q^{(k)}\|p_{A,H^{(k-1)}})$ (25)
$H^{(k)}$ $=\operatorname*{arg\,min}\limits_{H\in S^{d}_{\succeq 0}}D_{KL}(q^{(k)}\|p_{A^{(k)},H}).$ (26)
In iteration $k$, we start by using the SDE parameters estimated from the previous iteration to define a reference SDE $p_{A^{(k-1)},H^{(k-1)}}\in\mathcal{A}$. Using this reference SDE and the set of observed temporal marginals, we perform trajectory inference (24) (see middle plot of Figure 3). This problem is equivalent to a Schrödinger Bridge problem [84, 7], which seeks the path distribution $q^{(k)}\in\mathcal{D}$ that minimizes relative entropy to the reference SDE, while satisfying the marginal constraints (22). After trajectory inference, we perform a two-step maximum likelihood estimation, in order to estimate the drift and diffusion. This in turn updates the law of the reference SDE $p_{A^{(k)},H^{(k)}}\in\mathcal{A}$ for trajectory inference in the next iteration.
We note that this procedure can easily be adapted to a general additive noise SDE (1), provided that a corresponding maximum likelihood estimator is used for the nonlinear drift function $b^{(k)}(X_{t})$, in place of $A^{(k)}$, in the drift update step (25). We show in Corollary 8.19 of the Appendix that the structure of the closed form diffusion MLE is unchanged for general additive noise SDEs.
Figure 3: Visualization of our iterative parameter estimation algorithm APPEX. Given observed temporal marginals from an underlying SDE (left), APPEX alternates between trajectory inference (middle) and MLE parameter estimation (right) in order to find the SDE parameters that best represent the temporal marginal observations.
Previous works [93, 74, 84] implement similar iterative schemes, which alternate between trajectory inference and MLE parameter estimation. In contrast to our method, each of these works assumes that the diffusion of the process is known, and hence invariant across iterations. One theoretical reason for this assumption is to ensure finite KL divergence between the estimated trajectories and estimated SDEs for all iterations [84]. Given continuously observed marginals of a $d$-dimensional process, the KL divergence between two laws on paths $q,p$ is taken over the path space $\Omega=C([0,T],\mathbb{R}^{d})$, such that
$D_{KL}(q\|p)=\int_{\Omega}\log\left(\frac{\mathrm{d}q}{\mathrm{d}p}(\omega)\right)\mathrm{d}q(\omega).$
Thus, $D_{KL}(q\|p)$ will only be finite between two laws on paths if we can define the Radon-Nikodym derivative $\frac{\mathrm{d}q}{\mathrm{d}p}$ over path space $\Omega$. By Girsanov’s theorem, this is only ensured when both processes share the same diffusion [74, 84]. However, if we only consider measurements from a finite number of observed marginals, the KL divergence of the discretized processes $q^{N},p^{N}$ over the path space projected on the $N$ measurement times, $\Sigma_{\mathcal{D}}=C(\{t_{i}\}_{i=0}^{N-1},\mathbb{R}^{d})$, can be decomposed as follows, see [6]:
$D_{KL}(q^{N}\|p^{N})=\sum_{i=0}^{N-2}D_{KL}(q_{i,i+1}\|p_{i,i+1})-\sum_{i=1}^{N-2}D_{KL}(q_{i}\|p_{i}),$ (27)
where $p_{i,i+1},q_{i,i+1}$ are the joint probability measures of $q,p$, restricted to times $t_{i}$ and $t_{i+1}$, and similarly $q_{i},p_{i}$ are the marginals of $q,p$ at time $t_{i}$. We note that $D_{KL}(q_{i}\|p_{i})\leq D_{KL}(q_{i,i+1}\|p_{i,i+1})$ holds by the data-processing inequality. Hence, $D_{KL}(q^{N}\|p^{N})$ will be finite as long as $D_{KL}(q_{i,i+1}\|p_{i,i+1})=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\log\left(\frac{\mathrm{d}q_{i,i+1}}{\mathrm{d}p_{i,i+1}}(x)\right)q_{i,i+1}(x)\,\mathrm{d}x<\infty$ for each $i=0,\ldots,N-1$. In particular, if $p$ and $q$ are the laws of two different drift-diffusion SDEs with distinct non-degenerate diffusions, then the KL divergence $D_{KL}(q^{N}\|p^{N})$ over the discretized path space $\Sigma_{\mathcal{D}}$ will be finite. Since we consider a setting with a finite number of temporal marginal observations in this work, we may consider laws on paths discretized over these observations. This allows us to both consider diffusions which are not known in advance, and improve our diffusion estimates with respect to KL divergence after each iteration.
5.1 Trajectory inference via anisotropic entropy-regularized optimal transport
To begin each iteration of APPEX, we need to infer trajectories, given the set of temporal marginals $\{p_{i}\}_{i=0}^{N-1}$, and a reference SDE, $p_{A^{(k-1),H^{k-1)}}}\in\mathcal{A}$. As previously discussed, this is equivalent to the Schrödinger Bridge problem (24), which may be solved via entropy-regularized optimal transport (EOT). Given two marginals $\mu,\nu$, the solution to the standard EOT problem of transporting probability measure $\mu$ to $\nu$ with entropic regularization $\epsilon>0$ and cost $c(x,y)=\|y-x\|^{2}/2$ is
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi(\mu,\nu)}\int c(x,y)\,\mathrm{d}\pi(x,y)+\epsilon^{2}D_{KL}(\pi\|\mu\otimes\nu)=\operatorname*{arg\,min}_{\pi\in\Pi(\mu,\nu)}\int\frac{\|y-x\|^{2}}{2\epsilon^{2}}\,\mathrm{d}\pi(x,y)+D_{KL}(\pi\|\mu\otimes\nu).$ (28)
Furthermore, it is easy to see from the dual formulation (see (53) in the Appendix) that $\pi^{*}\in\Pi(\mu,\nu)$ is the joint distribution with marginals $\mu,\nu$, which minimizes relative entropy with respect to the reference measure $d\mathcal{K}(x,y)=K(x,y)d\mu(x)d\nu(y)$ with density $K(x,y)\propto e^{\frac{c(x,y)}{\epsilon^{2}}}\propto e^{\frac{-\|y-x\|^{2}}{2\epsilon^{2}}}$ [66, 93]:
$\pi^{*}=\text{Proj}_{\Pi(\mu,\nu)}^{KL}(\mathcal{K})=\operatorname*{arg\,min}_{\pi\in\Pi(\mu,\nu)}D_{KL}(\pi\|\mathcal{K}).$ (29)
Remark 5.1 (Applying EOT for SDE trajectory inference).
In the standard $\epsilon$-regularized OT problem (28), $K\sim\mathcal{N}(x,\epsilon^{2}\mathrm{I}_{d})$ is an isotropic Gaussian kernel. As noted in [47], this implies that entropy regularized OT can be leveraged for trajectory inference from observed marginals, given a reference SDE. For example, to find the discretized law on paths $\pi^{*}\in\Pi(\mu,\nu)$ satisfying $P(X_{t}=\mu,X_{t+\Delta t}=\nu)$, which minimizes relative entropy to the law of a pure diffusion process $\mathrm{d}X_{t}=\sigma\mathrm{d}W_{t}$, one should set the entropic regularization $\epsilon^{2}=\sigma^{2}\Delta t$. Indeed, this would correspond to minimizing the KL divergence to $K(x,y)=e^{-\frac{\|y-x\|^{2}}{2\sigma^{2}\Delta t}}\sim\mathcal{N}(x,\sigma^{2}\Delta t)$, which is the transition kernel of the reference SDE. Similarly, as done in [93], one can perform trajectory inference given an Ornstein-Uhlenbeck reference SDE $\mathrm{d}X_{t}=-AX_{t}\mathrm{d}t+\sigma\mathrm{d}W_{t}$ by approximating the transition distribution $X_{t+\Delta t|t}$ via $K(x,y)=e^{-\frac{\|y-e^{A\Delta t}x\|^{2}}{2\sigma^{2}\Delta t}}$. This would correspond to reweighting the squared Euclidean cost with the drift matrix $A$, such that $c(x,y)=\|y-e^{A\Delta t}x\|/2$, and applying standard entropy regularized OT with $\epsilon^{2}=\sigma^{2}\Delta t$.
However, the standard EOT problem (28) only considers a scalar regularization parameter $\epsilon$, which means that it can only directly model reference SDEs with isotropic diffusion. To generalize trajectory inference for SDEs with anisotropic diffusion, we formalize “anisotropic entropy-regularized optimal transport” (AEOT), by considering transition kernels $K$ with custom mean and covariance, e.g. parameterizing $K$ with $A$ and $H=GG^{\top}$ rather than the scalar $\epsilon^{2}>0$. Indeed, we can consider more general Gaussian transition kernels, $K_{\theta}(x,y)=\exp({\frac{(y-\mu(x,\theta))^{\top}\Sigma(\theta)^{-1}(y-\mu(x,\theta))}{2}})$, in order to model a transition $y|x\sim\mathcal{N}(\mu(x,\theta),\Sigma(\theta))$, whose covariance $\Sigma(\theta)$ is possibly anisotropic. Under this formulation, the cost is given by the inner product $c_{\theta}(x,y)=\frac{(y-\mu(x,\theta))^{\top}\Sigma(\theta)^{-1}(y-\mu(x,\theta))}{2}$. Trajectory inference would amount to solving the AEOT problem:
$\pi^{*}=\operatorname*{arg\,min}_{\pi\in\Pi(\mu,\nu)}\int\frac{(y-\mu(x,\theta))^{\top}\Sigma(\theta)^{-1}(y-\mu(x,\theta))}{2}\,\mathrm{d}\pi(x,y)+D_{KL}(\pi\|\mu\otimes\nu),$ (30)
where the entropic regularization is captured in the matrix $\Sigma(\theta)$ and function $\mu(x,\theta)$. Just as in the standard EOT problem, the AEOT solution $\pi^{*}$ can be obtained by Sinkhorn’s algorithm.
Proposition 8.18 in the Appendix shows that the solution to the trajectory inference step (24) is the joint distribution given by the Markov concatenation of couplings (see Definition 8.17 in the Appendix)
$\pi^{*}=\pi_{0,1}\circ\ldots\circ\pi_{N-2,N-1},$
where $\pi_{i,i+1}$ is the AEOT solution (30) with marginals $\mu=p_{i},\nu=p_{i+1}$ and transition kernel
$K_{A,H}^{i}(x,y)\propto\exp(-\frac{1}{2}(y-e^{A(t_{i+1}-t_{i})}x)^{\top}(\Sigma_{i})^{-1}(y-e^{A(t_{i+1}-t_{i})}x)$ (31)
such that $\Sigma_{i}=\int_{t_{i}}^{t_{i+1}}e^{A(t_{i+1}-s)}He^{A^{\top}(t_{i+1}-s)}\,\mathrm{d}s$. The reference transition $X_{t_{i+1}}|X_{t_{i}}\sim\mathcal{N}(e^{A(t_{i+1}-t_{i})}X_{t},\Sigma_{t})$ can be estimated via the first order approximation $\mathcal{N}(X_{t}+AX_{t}(t_{i+1}-t_{i}),H(t_{i+1}-t_{i}))$.
Remark 5.2 (Application to empirical marginals).
Proposition 8.18 applies in the case where one observes the exact temporal marginals $p_{i}$. However, in practice we observe empirical measures $\hat{p}_{i}$ which, in the limit of infinite data, converge almost surely in distribution to $p_{i}$. It is for these empirical measures $\hat{p}_{i}$ which we actually compute the estimated couplings $\hat{\pi}_{i,i+1}$ in practice. This corresponds to numerically estimating the AEOT solution
$\hat{\pi}_{i,i+1}=\operatorname*{arg\,min}_{\pi\in\Pi(\hat{p}_{i},\hat{p}_{i+1})}\int\frac{(y-\mu(x,\theta))^{\top}\Sigma(\theta)^{-1}(y-\mu(x,\theta))}{2}\,\mathrm{d}\pi(x,y)+D_{KL}(\pi\|\hat{p}_{i}\otimes\hat{p}_{i+1}).$ (32)
By the main theorem in [31], it holds that $\hat{\pi}_{i,i+1}$ converges in distribution almost surely to the minimizer of (30) with $\mu=p_{i}$ and $\nu=p_{i+1}$ as the number of samples goes to infinity for each time $i$ and $i+1$. Combining this with the previous proposition, we see that estimating (32) for each pair of times $i$ and $i+1$ is asymptotically equivalent to the KL minimization step in (24).
We also note that (32) is practically solved using Sinkhorn’s algorithm with empirical marginals $\hat{p_{i}},\hat{p}_{i+1}$ and the discretized transition kernel $K_{\theta}$ of the reference SDE. For example, if we have $M$ samples per empirical marginal, then $\hat{p}_{i}\sim\text{Unif}(x_{t_{i}}^{(j)}:j=1,\ldots M)$ and $\hat{p}_{i+1}\sim\text{Unif}(x_{t_{i+1}}^{(j)}:j=1,\ldots M)$ would both be discrete uniform distributions over their samples, and $K_{\theta}\in\mathbb{R}^{M\times M}$ would be a square matrix, such that entry $K_{\theta_{j,k}}$ is obtained by applying the given transition kernel on the data points $x_{t_{i}}^{(j)}$ and $x_{t_{i+1}}^{(k)}$.
5.2 Parameter estimation via MLE
To optimize objectives (25) and (26) for each iteration of APPEX, we require maximum likelihood estimators for the SDE parameters in the setting of multiple observed trajectories from $[0,T]$. We derive closed-form maximum likelihood estimators for the linear additive noise SDE from Equation 2, given multiple trajectories. In the context of iteration $k$ of APPEX, these are the trajectories sampled from the law on paths $q^{(k)}$ obtained from the trajectory inference step.
Proposition 5.3 (MLE estimators for drift and diffusion of SDE (2) from multiple trajectories).
Given $M$ trajectories over $N$ equally spaced times: $\{X_{i\Delta t}^{(j)}:i\in 0,...,N-1,\text{ }j\in 0,...,M-1\}$ sampled from the linear additive noise SDE (2), the maximum likelihood solution for linear drift is approximated by
$\hat{A}$ $=\frac{1}{\Delta t}\left(\sum_{i=0}^{N-2}\sum_{j=0}^{M-1}(\Delta X_{i}^{(j)})X_{i}^{{(j)}^{\top}}\right)\left(\sum_{i=0}^{N-2}\sum_{j=0}^{M-1}X_{i}^{(j)}X_{i}^{{(j)}^{\top}}\right)^{-1}$ (33)
$\overset{\begin{subarray}{c}M\to\infty\\[2.0pt] \end{subarray}}{\longrightarrow}\frac{1}{\Delta t}\left(\sum_{i=0}^{N-2}\mathbb{E}_{p_{i,i+1}}[(\Delta X_{i})X_{i}^{\top}]\right)\left(\sum_{i=0}^{N-1}\mathbb{E}_{p_{i}}[X_{i}X_{i}^{\top}]\right)^{-1},$ (34)
where $p_{i,i+_{1}}$ is the joint measure over $X_{t_{i}}$ and $X_{t_{i+1}}$. Similarly, the maximum likelihood solution for diffusion is approximated by
$\hat{H}$ $=\frac{1}{MT}\sum_{i=0}^{N-2}\sum_{j=0}^{M-1}\left(\Delta X_{i}^{(j)}-AX_{{i}}^{(j)}\Delta t)(\Delta X_{i}^{(j)}-AX_{{i}}^{(j)}\Delta t)^{\top}\right)$ (35)
$\overset{\begin{subarray}{c}M\to\infty\\[2.0pt] \end{subarray}}{\longrightarrow}\frac{1}{T}\sum_{i=0}^{N-2}\mathbb{E}_{p_{i,i+1}}[\left(\Delta X_{i}-AX_{i}\Delta t\right)\left(\Delta X_{i}-AX_{i}\Delta t\right)^{\top}]$ (36)
Proof.
See Section 8.6 of the Appendix. We note that estimators $\hat{A}$ and $\hat{H}$ were derived using the discretized transition kernel, $X_{i+1}|X_{i}\sim\mathcal{N}(X_{i}+AX_{i}\Delta t,Hdt)$. We derive the maximum likelihood estimators with the exact transition kernel $X_{i+1}|X_{i}\sim\mathcal{N}(e^{A\Delta t}X_{i},H\Delta t)$ in the one dimensional case in Section 8.6. ∎
Remark 5.4.
Note that the MLE estimator for drift $A$ does not depend on the diffusion $H$, but the MLE estimator for $H$ depends on $A$. Therefore, we estimate drift first in each iteration of APPEX.
5.3 The APPEX algorithm
APPEX has two main subprocedures. We first perform a trajectory inference step (24), via anisotropic entropy-regularized optimal transport (AEOT), our adaptation of EOT for general additive noise SDEs. This step is outlined in line $6$ of Algorithm 1. We then perform two maximum likelihood parameter estimation steps (25)-(26), where we have derived approximate MLE solutions for linear drift and additive diffusion in Proposition 5.3. These steps are outlined in lines $9$ and $10$ of Algorithm 1. Because these subprocedures are optimal with respect to minimizing KL divergence, we can show that APPEX’s estimates will improve with each iteration.
Lemma 5.5.
Suppose that $A$ and $H$ are the true drift and diffusion parameters of a linear additive noise SDE with temporal marginals $p_{0},...,p_{N-1}$. Let $q^{(k)}\in\mathcal{D}$ be the law of the estimated trajectories at iteration $k$ of APPEX, and $p_{A^{k},H^{k}}\in\mathcal{A}$ be the law of the law of the estimated linear additive noise SDE, such that $A^{0}\in\mathbb{R}^{d\times d}$ and $H^{0}\in S^{d}_{\succeq 0}$. Then, the relative entropy between the estimated law of the trajectories and the law of the estimated linear additive noise SDE is decreasing with each iteration:
$D_{KL}(q^{(k+1)}\|p_{A^{k+1},H^{k+1}})\leq D_{KL}(q^{(k)}\|p_{A^{k},H^{k}})\ \ \forall k\geq 0.$
Furthermore, if $p_{0}$ is not auto-rotationally invariant, then as the number of observed marginals $N\to\infty$, the relative entropy between the estimated law of the trajectories and the law of the estimated SDE is uniquely minimized by the underlying SDE $p_{A,H}$:
$\inf\limits_{q\in\mathcal{D},p\in\mathcal{A}}D_{KL}(q\|p)=0\iff q=p=p_{A,H}$
Proof.
Proposition 8.18 shows that $q^{(k+1)}\in\mathcal{D}$ minimizes relative entropy to $p_{A^{k},H^{k}}$. Similarly, any MLE solution $p_{A^{(k+1)},H^{(k+1)}}$ minimizes relative entropy to $q^{(k+1)}$. In particular, Proposition 5.3 approximates the MLE parameters $A^{k+1}\in\mathbb{R}^{d\times d}$ and $H^{(k+1)}\in S^{d}_{\succeq 0}$. By construction of the iterative scheme, we have that
$D_{KL}(q^{(k+1)}\|p_{A^{(k+1)},H^{(k+1)}})\leq D_{KL}(q^{(k+1)}\|p_{A^{(k+1)},H^{(k)}})\leq D_{KL}(q^{((k+1))}\|p_{A^{(k)},H^{(k)}})\leq D_{KL}(q^{(k)}\|p_{A^{(k)},H^{(k)}}),$
which proves the first claim of the lemma. Furthermore, if $X_{0}$ is not auto-rotationally invariant, then by our identifiability result (Theorem 4.3), as we increase the number of temporal marginals $N\to\infty$, $p_{A,H}$ is the law of the unique linear additive noise SDE, which obey the marginal constraints $p_{0},\ldots p_{N-1}$. Equivalently, we have $\mathcal{A}\cap\mathcal{D}=\{p_{A,H}\}$ and
$\inf\limits_{q\in\mathcal{D},p\in\mathcal{A}}D_{KL}(q\|p)=0\iff q=p=p_{A,H},$
Since the relative entropy $D_{KL}(q^{(k)}\|p_{A^{(k)},H^{(k)}})$ is decreasing for successive iterations $k$, this implies that the drift and diffusion estimates $A^{(k)},H^{(k)}$ increasingly fit the data with each iteration. In this sense, our method is always approaching the true solution, which is unique, provided that it is given an appropriate initial distribution and sufficient observations.
Algorithm 1 Parameter estimation for an additive noise SDE from temporal marginals with APPEX
1:Input: Observed marginals $\hat{p}_{i}$, $i=0,\ldots,N-1$, number of iterations $K$, $\Delta t$
2:Result: Estimated drift function $\hat{b}$ and additive noise $\hat{H}$
3:Initialize: $\hat{b}\leftarrow 0$, $H\leftarrow\sigma^{2}\mathrm{I}_{d}$, $k\leftarrow 0$
4:while $k<K$ do
5:     for $i=1,\ldots,N$ do
6:         $\hat{\pi}_{i,i+1}\leftarrow\text{Anisotropic-Entropy-Regularized-Optimal-Transport}(\hat{b},\hat{H},\hat{p}_{i-1},\hat{p}_{i},\Delta t)$
7:     end for
[…]
10:     $\hat{H}\leftarrow\text{MLEfit}(\text{Sample-Trajectories},\hat{b})$
11:     $k\leftarrow k+1$
12:end while
[…]
Remark 5.6 (Application to causal discovery).
As proven in Lemma 4.5 and visualized in Figure 1, the causal graph can be derived from the SDE parameters. For nonlinear drift $b(X_{t})$, we include the edge $i\to j$ if the $j$th component of the drift, $b_{j}$, is a function of $X^{(i)}$. For linear drift one may simply examine $A_{j,i}$. For multi-edges, one can check $H_{i,j}$ to determine whether there is an unobserved confounder causing $X^{(i)}$ and $X^{(j)}$. We therefore include a causal discovery step in line $13$ of Algorithm 1. We acknowledge that many causal algorithms and independence tests are applicable, and leave the implementation as general as possible.
6 Experiments
In this section, we evaluate the performance of our parameter estimation method APPEX on simulated data. We first demonstrate that the non-identifiable examples from Section 3 can be correctly identified by APPEX if the initial marginal $X_{0}$ is not auto-rotationally invariant. We then show that combining our initialization with our parameter estimation method successfully learns arbitrary dynamical systems modelled with a linear additive noise SDE by testing it on randomly generated SDEs across a range of dimensions. We conclude this section with several causal graph recovery experiments. An empirical experiment for consistency is also provided in Appendix 9.1, which shows that over a broad class of random experiments, APPEX’s estimates for both drift and diffusion converge to the true SDE parameters.
6.1 Experimental setup
For each experiment, we simulate data for linear additive noise SDEs using Euler-Maruyama discretization with $dt_{EM}=0.01$, such that we generate $M=500$ trajectories, each observed across $100$ time steps. Each $d$-dimensional trajectory is initialized with $p_{0}\sim\text{Unif}\{x_{i}\}_{i=1}^{d}$ where $\{x_{i}\}_{i=1}^{d}$ are randomly sampled linearly independent vectors with each entry having a magnitude between $2$ and $10$. It follows from Proposition 8.11 and Theorem 4.3, that this initial distribution ensures that parameter estimation is feasible.
To model the setting where we only observe temporal marginals, we subsample at the rate $dt=0.05$ to produce $N=20$ marginals with $M=500$ observations per time. To perform parameter estimation, we use $30$ iterations of our APPEX algorithm, such that the initial reference SDE is an isotropic Brownian motion $\mathrm{d}X_{t}=\sigma\mathrm{d}W_{t}$, even when the true process has anisotropic noise. To model the realistic setting where diffusion is not precisely known, we assume that the initial guess of the diffusion’s trace is within an order of magnitude of the ground truth trace $\text{tr}(H)$. Specifically, we initialize $A^{(0)}=0,H^{(0)}=\sigma^{2}\mathrm{I}_{d}\ s.t.\ \sigma^{2}\sim\text{tr}(H)10^{\text{Unif}(-1,1)}$. Because the first reference SDE is $\sigma\mathrm{d}W_{t}$, we note that the first iteration of APPEX is equivalent to inferring trajectories using the Waddington-OT (WOT) method from [78], based on standard regularized OT (28) with $\epsilon^{2}=\sigma^{2}\Delta t$, followed by MLE parameter estimation. APPEX and WOT are distinguished by the fact that APPEX benefits from further iterations, and allows the reference SDE to have anisotropic diffusion and non-zero drift. For each iteration of APPEX, we use Sinkhorn’s algorithm to solve the anisotropic entropy-regularized optimal transport problem for trajectory inference, and we use the closed form MLE solutions from Proposition (5.3). Due to time complexity and numerical stability, we use linearized discretizations for the Gaussian transition kernels (unless $d=1$).
6.2 Revisiting previously non-identifiable SDEs
We perform parameter estimation on the three pairs of SDEs from Section 4. $10$ replicates of each experiment were performed, such that each replicate featured data from different valid $p_{0}$ initializations and different initial diffusivities $\sigma^{2}$. To measure performance, we track the mean absolute error (MAE), plotted in Figure 4, between the true drift/diffusion parameters and their estimates.
(a) $\mathrm{d}X_{t}=-X_{t}\>\mathrm{d}t+\mathrm{d}W_{t}$ from Example 3.1
(b) $\mathrm{d}X_{t}=-10X_{t}\>\mathrm{d}t+\sqrt{10}\mathrm{d}W_{t}$ from Ex.  3.1
(c) $\mathrm{d}X_{t}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\mathrm{d}W_{t}$ from Ex.  3.2
(d) $\mathrm{d}X_{t}=\begin{bmatrix}0&1\\ -1&-0\end{bmatrix}X_{t}\mathrm{d}t+\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\mathrm{d}W_{t}$ from Ex.  3.2
(e) $\mathrm{d}X_{t}=\begin{bmatrix}1&2\\ 1&0\end{bmatrix}X_{t}\mathrm{d}t+\begin{bmatrix}1&2\\ -1&-2\end{bmatrix}\mathrm{d}W_{t}$ from Ex.3.3
(f) $\mathrm{d}X_{t}=\begin{bmatrix}\frac{1}{3}&\frac{4}{3}\\ \frac{2}{3}&-\frac{1}{3}\end{bmatrix}X_{t}\mathrm{d}t+\begin{bmatrix}1&2\\ -1&-2\end{bmatrix}\mathrm{d}W_{t}$ from Ex.3.3
Figure 4: The mean absolute error for estimates of $A$ and $H$ using APPEX is shown per iteration for all three pairs of SDEs from Section (4)
The results demonstrate that in each example, APPEX is able to estimate both drift and diffusion by iteratively improving upon both estimates. In each experiment, we observe decreasing MAEs for both SDE parameters as the number of iterations increases. The worst performance was observed for the SDE $\mathrm{d}X_{t}=-10X_{t}\>\mathrm{d}t+\sqrt{10}\mathrm{d}W_{t}$, whose significantly higher noise scale made inference more difficult, particularly for the diffusion, which APPEX consistently estimated around $H=7$ rather than $H=10$. We note that even though degenerate diffusions can result in infinite KL divergence with respect to a misspecified reference SDE, we observe that APPEX can estimate degenerate diffusions, such as in Example 3.3.
6.3 Higher dimensional random matrices
We now test APPEX on a broad range of higher dimensional random linear additive noise SDEs. For each dimension $d=3,\ldots,10$, we generate $10$ random SDEs. To create each drift matrix $A$, we randomly sample each of its $d^{2}$ entries from $\text{Unif}(-5,5)$. To ensure that the values do not diverge, we verify that the maximal real part of the eigenvalues of $A$ is less than $1$. This allows us to consider process beyond OU processes, while obeying reasonable growth conditions in practice. To create each diffusion $G$, we randomly initialize each of its $d^{2}$ entries via $\text{Unif}(-1,1)$ and then set $H=GG^{\top}$. For numerical stability, we perform Sinkhorn on the logarthmic scale for this experiment.
Table 1: Mean absolute error (MAE) of estimated drift and diffusion for dimensions 3-10. Our method (APPEX) consistently outperforms (bold) the previous method (WOT).
Dimension A Estimation GGT Estimation
WOT APPEX WOT APPEX
3 $0.351\pm 0.04$ $\mathbf{0.237\pm 0.04}$ $0.793\pm 0.205$ $\mathbf{0.147\pm 0.030}$
4 $0.730\pm 0.067$ $\mathbf{0.328\pm 0.041}$ $1.549\pm 0.439$ $\mathbf{0.415\pm 0.070}$
5 $0.912\pm 0.060$ $\mathbf{0.602\pm 0.195}$ $2.174\pm 0.702$ $\mathbf{0.362\pm 0.039}$
6 $1.43\pm 0.170$ $\mathbf{0.358\pm 0.020}$ $18.010\pm 8.636$ $\mathbf{0.256\pm 0.046}$
7 $1.480\pm 0.132$ $\mathbf{0.360\pm 0.015}$ $6.807\pm 1.724$ $\mathbf{0.345\pm 0.037}$
8 $1.862\pm 0.137$ $\mathbf{0.460\pm 0.015}$ $5.472\pm 1.266$ $\mathbf{0.359\pm 0.019}$
9 $1.803\pm 0.222$ $\mathbf{0.487\pm 0.016}$ $8.134\pm 3.024$ $\mathbf{0.454\pm 0.122}$
10 $1.670\pm 0.241$ $\mathbf{0.439\pm 0.019}$ $35.122\pm 28.529$ $\mathbf{0.317\pm 0.025}$
Table 2: Correlation between estimated and true drift and diffusion for dimensions 3-10. Our method (APPEX) consistently outperforms (bold) the previous method (WOT).
Dimension A Estimation GGT Estimation
WOT APPEX WOT APPEX
3 $0.996\pm 0.001$ $\mathbf{0.998\pm 0.001}$ $0.837\pm 0.048$ $\mathbf{0.985\pm 0.005}$
4 $0.943\pm 0.015$ $\mathbf{0.987\pm 0.005}$ $0.729\pm 0.039$ $\mathbf{0.865\pm 0.031}$
5 $0.921\pm 0.016$ $\mathbf{0.952\pm 0.030}$ $0.728\pm 0.040$ $\mathbf{0.909\pm 0.018}$
6 $0.794\pm 0.040$ $\mathbf{0.986\pm 0.001}$ $0.530\pm 0.056$ $\mathbf{0.961\pm 0.007}$
7 $0.792\pm 0.029$ $\mathbf{0.988\pm 0.001}$ $0.595\pm 0.037$ $\mathbf{0.946\pm 0.012}$
8 $0.699\pm 0.035$ $\mathbf{0.981\pm 0.002}$ $0.611\pm 0.042$ $\mathbf{0.949\pm 0.006}$
9 $0.740\pm 0.033$ $\mathbf{0.978\pm 0.002}$ $0.615\pm 0.025$ $\mathbf{0.919\pm 0.033}$
10 $0.760\pm 0.049$ $\mathbf{0.983\pm 0.001}$ $0.641\pm 0.041$ $\mathbf{0.960\pm 0.006}$
(a) $d=3$
(b) $d=10$
Figure 5: The correlation between the estimated and true SDE parameters is plotted per iteration across $10$ random linear additive noise SDEs for dimensions $3$ and $10$
The results demonstrate that APPEX continues to estimate both SDE parameters robustly across all settings. Importantly, APPEX can handle arbitrary additive noise structures $H=GG^{\top}$, as evidenced by high correlations to the true diffusion and low MAE. This is significant because previous literature has focused on the setting of isotropic noise, but in practice, we expect noise structures to have off-diagonal entries due to shared noise from unobserved confounders, as well as unequal noise along the main diagonal. In contrast, although WOT estimates the drift somewhat decently, it is unable to estimate the diffusion accurately, particularly in higher dimensions, since it is constrained to isotropic noise in its reference SDE. Figure (5) shows how APPEX is able to re-orient incorrect diffusion priors to closely match the true diffusion of the underlying SDE.
6.4 Causal discovery
We conclude with experiments that demonstrate APPEX’s ability to recover the causal graph of the underlying dynamic system. We first consider causal discovery from systems without latent confounders, and then consider causal discovery from systems with latent pairwise confounders.
We evaluate APPEX using the default experimental settings laid out in Section 6.1 across randomly generated SDEs of dimension $d=3,5,10$. In particular, we consider Erdös-Renyi graphs $G(d,p)$, such that each of the $d(d-1)$ possible simple directed edges are included with probability $p$. As in Section 6.3, we ensure that the maximal eigenvalue of the randomly generated ground truth drift matrix is at most $1$ to prevent unbounded growth. We will see that accurately estimating the diffusion is important for recovering the correct causal graph.
6.4.1 Causal discovery under causal sufficiency
Lemma 4.5 states that simple directed edges $e=i\to j\in E$ in the causal graph $\mathcal{G}=([d],E,\tilde{E})$ are characterized by the condition $A_{j,i}\neq 0$. To simulate ground truth edges, we follow standard convention for simulating data for causal discovery, by simulating edge weights $A_{j,i}$ uniformly via $\text{Unif}(-5,0.5)\cup\text{Unif}(0.5,5)$ [72, 71]. We then determine the presence of edge $i\to j$, if in the estimated drift, $\hat{A}_{j,i}>\epsilon$ (corresponding to a positive edge weight) or $\hat{A}_{j,i}<-\epsilon$ (corresponding to a negative edge weight). We choose our threshold $\epsilon=0.5$ according to the minimal edge weight magnitude from simulated construction.
To model the case where the observed system does not have any latent confounders, we set the diffusion $G$ to be zero for all entries outside its main diagonal. Thus, the only edges in the causal graph $\mathcal{G}$ are simple directed edges, attributed to the drift. Similarly to the experiment on higher dimensional random matrices, we set each of the $d$ diagonal entries of $G$ via $\text{Unif}(0,1)$.
We measure performance according to the structural Hamming distance. In particular, a system’s causal graph consists of directed edges $i\to j$ with positive weights $A_{j,i}>0$ ($\hat{A}_{j,i}>\epsilon$), directed edges $i\to j$ with negative weights $A_{j,i}<0$ ($\hat{A}_{j,i}<-\epsilon$), and absence of edges $A_{j,i}=0$ ($|\hat{A}_{j,i}|<\epsilon$). The structural Hamming distance adds $1$ for every instance in which an edge is misclassified, and is defined by
$d(\mathcal{G}(A),\mathcal{G}({\hat{A}}))=\sum_{(i,j)\in[d]\times[d]}\mathbf{1}\left\{\operatorname{sgn}(A_{j,i})\neq\operatorname{sgn}(\hat{A}_{j,i})\mathbf{1}_{|\hat{A}_{j,i}|>\epsilon}\right\}.$ (37)
The mean structural Hamming distances of causal graphs estimated by WOT and APPEX, across various dimensions $d\in\{3,5,10\}$ and random edge probabilities $p\in\{0.1,0.25,0.5\}$, are given in Table 3. We also plot the true vs. estimated graphs by WOT and APPEX for two SDEs in Figure 6.
While WOT often struggles to learn the causal graph, due to drift estimates compensating for mispecified diffusion [36], APPEX consistently recovers the causal graph, including cycles and v-structures. For example, the first row in Figure 6 demonstrates that APPEX recovers the system’s negative feedback structure, while WOT introduces additional cycles between variables to make sense of the data with poorly estimated diffusion. This highlights the fact that one must estimate diffusion accurately in order to recover an accurate causal representation of the system from the drift.
Figure 6: The true and estimated causal graphs by WOT and APPEX for two random SDEs of dimensions $d=5,10$ are illustrated. Red edges $i\to j$ represent negative edge weights, such that $X^{(i)}$ negatively regulates $X^{(j)}$. Green edges $i\to j$ represent positive edge weights, such that $X^{(i)}$ positively regulates $X^{(j)}$.
Table 3: Average Structural Hamming Distance (lower is better) with varying dimensions and random edge probabilities $p$. Our method (APPEX) consistently outperforms (bold) the previous method (WOT).
Dimension $p=0.1$ $p=0.25$ $p=0.5$
WOT APPEX WOT APPEX WOT APPEX
3 $0.40\pm 0.40$ $\mathbf{0.00\pm 0.00}$ $1.80\pm 0.76$ $\mathbf{0.30\pm 0.30}$ $1.40\pm 0.65$ $\mathbf{0.00\pm 0.00}$
5 $3.80\pm 2.13$ $\mathbf{0.10\pm 0.10}$ $3.56\pm 1.70$ $\mathbf{0.00\pm 0.00}$ $4.40\pm 1.14$ $\mathbf{0.60\pm 0.22}$
10 $38.30\pm 8.53$ $\mathbf{0.50\pm 0.40}$ $48.60\pm 6.55$ $\mathbf{0.80\pm 0.39}$ $37.4\pm 4.66$ $\mathbf{3.30\pm 0.70}$
6.4.2 Causal discovery with latent confounders
Lemma 4.5 also states that pairwise latent confounders, represented by bi-directed edges $\tilde{e}=i\leftrightarrow j$ in the causal graph $\mathcal{G}=([d],E,\tilde{E})$ are characterized by the condition $H_{i,j}=H_{j,i}\neq 0$. Since we would also like to consider cycles formed by sets of simple directed edges, we consider the augmented causal $\bar{\mathcal{G}}=(V,E)$ to more clearly model pairwise latent confounders, as done in [11]. Specifically, $\tilde{e}=i\leftrightarrow j\in\tilde{E}$ would correspond to the v-structure $i\leftarrow U_{i,j}\rightarrow j$ in the augmented graph $\bar{\mathcal{G}}$, where $U_{i,j}$ is the unobserved pairwise confounder of $X^{(i)}$ and $X^{(j)}$.
To model the case where the observed system features pairwise latent confounders, we randomly select a subset of the columns of the diffusion $G$ to feature precisely two nonzero entries, in two randomly chosen rows $i,j$. For simplicity, we initialize these entries $G_{i,k}=G_{j,k}=1\implies H_{i,j}=G_{i}\cdot G_{j}\geq 1$. We then determine the presence of the v-structure $i\leftarrow U_{i,j}\rightarrow j$ if in the estimated observational diffusion, $|\hat{H}_{i,j}|=|\hat{H}_{j,i}|>\epsilon$. We again pick our threshold to be $\epsilon=0.5$.
We again measure performance according to the structural Hamming distance. For interpretability, we break the distance into two parts. We consider the Hamming distance based on the simple edges between observed variables, defined in (37), as well as the Hamming distance based on latent pairwise confounders, which adds $1$ for every misclassified v-structure in the augmented graph:
$d(\bar{\mathcal{G}}(H),\bar{\mathcal{G}}(\hat{H})=\sum_{(i,j)\in[d]\times[d]:i\neq j}\mathbf{1}\left\{(|H_{i,j}|>\epsilon\cap|\hat{H}_{i,j}|\leq\epsilon)\cup(|H_{i,j}|\leq\epsilon\cap|\hat{H}_{i,j}|>\epsilon)\right\}.$ (38)
We evaluate APPEX and WOT across random SDEs of dimensions $d=3,5,10$ with random edge probability $p=0.25$ for simple edges. The number of pairwise latent confounders is chosen uniformly from $\{1,\dots,\left\lfloor\frac{2d}{3}\right\rfloor\}$. The results are summarized in Table 4. We also plot the true vs. estimated graphs by WOT and APPEX for two SDEs in Figure 7. We see that APPEX is able to recover both the simple edges and the latent pairwise confounders with high accuracy across dimensions. In constrast, WOT struggles to detect both, especially for higher dimensions. We note that in this setting with anisotropic diffusion, WOT’s initial diffusion reference is not only mispecified according to diagonal scaling, but further mispecifies the nonzero entries.
Table 4: Average Structural Hamming Distance for simple edges, and latent confounders (lower is better) with varying dimensions and random edge probability $p=0.25$. Our method (APPEX) consistently outperforms (bold) the previous method (WOT).
Dimension Hamming distance for simple edges Hamming distance for latent confounders
WOT APPEX WOT APPEX
3 $1.20\pm 0.51$ $\mathbf{0.7\pm 0.40}$ $1.00\pm 0.30$ $\mathbf{0.00\pm 0.00}$
5 $4.80\pm 1.47$ $\mathbf{1.60\pm 0.87}$ $3.50\pm 0.64$ $\mathbf{0.10\pm 0.10}$
10 $38.00\pm 6.30$ $\mathbf{3.00\pm 0.77}$ $23.80\pm 3.86$ $\mathbf{2.90\pm 1.39}$
Figure 7: The true and estimated augmented causal graphs by WOT and APPEX for two random SDEs of dimensions $d=3,5$ are illustrated. Red edges $i\to j$ represent negative edge weights, such that $X^{(i)}$ negatively regulates $X^{(j)}$. Green edges $i\to j$ represent positive edge weights, such that $X^{(i)}$ positively regulates $X^{(j)}$. Dashed edges represent the effect of a latent confounder $U_{i,j}$ on observed variables $i,j$.
7 Discussion
This work presents the first comprehensive approach for estimating both the drift and diffusion of SDEs from temporal marginals. We achieve this by proving that arbitrary time-homogeneous linear additive noise SDEs are identifiable from temporal marginals, provided that the initial distribution is not auto-rotationally invariant. This result eliminates the need for prior assumptions that drift or diffusion are known, as well as restrictive model simplifications, namely irrotational drift or isotropic diffusion, which were previously required to resolve non-identifiability. Additionally, we show that drift and diffusion identify the system’s causal graph. In sum, our theoretical contributions demonstrate that it is possible to precisely learn a process’s underlying dynamics—even in the presence of complex interactions, such as negative feedback loops and repressilator dynamics, or latent confounders—solely from temporal marginals. These insights have the potential to advance methods for trajectory inference and gene regulatory network (GRN) inference in single-cell biology.
In order to practically learn a process’ underlying dynamics from temporal marginals, we introduce the first method to simultaneously learn drift and diffusion from temporal marginals. To estimate arbitrary drift and additive noise, we apply anisotropic entropy-regularized optimal transport for anisotropic diffusions, and we pair this with maximum likelihood estimates of drift and diffusion. Our parameter estimation method, APPEX, then iterates towards the optimal parameter set by alternating between trajectory inference and parameter estimation. Given a non-auto-rotationally invariant initial distribution $X_{0}$, we show that APPEX performs effectively across extensive experiments on simulated data from arbitrary linear additive noise SDEs. In particular, APPEX is able to recover the process’ drift, diffusion, and causal graph, in contrast to WOT, which is currently one of the most popular methods.
In terms of practical significance, we remark that non-auto-rotational invariance is satisfied by a large class of distributions. Hence, one can easily check the initial distribution, or perform an intervention if needed, to guarantee system identifiability from data generated by a linear additive noise SDE. Furthermore, we have proven the asymptotic optimality of each of APPEX’s subprocedures (Lemma 5.5): given a finite set of the process’ true marginals, APPEX improves its estimates with each iteration. Our experiments in the finite data setting ($500$ samples per temporal marginal) further motivate its practical use. Unlike WOT, APPEX accurately estimates drift, diffusion, and causal structure across hundreds of randomly generated SDEs between dimensions $3$ to $10$. APPEX therefore offers significant inference capability and flexibility compared to previous methods, since it can fully learn the dynamic system, without relying on prior knowledge. Even if a user is only interested in one parameter, APPEX offers the ability to estimate it, regardless of how the other parameter is initialized. This is particularly significant given that a misspecified diffusion typically leads to a misidentified drift, and vice-versa.
7.1 Limitations and future work
Although this paper presents several breakthroughs for the identification of SDEs from marginals, we acknowledge methodological limitations and opportunities for future work.
First, while we have proved that identifiability holds when the initial distribution $X_{0}$ is not auto-rotationally invariant, we have not yet studied the relative effectiveness of different initial distributions for parameter estimation. This could be an interesting area for further exploration, since the initial distribution may, at times, be controllable.
Second, our identifiability result is restricted to time homogeneous linear additive noise SDEs. Although this is a popular model for real-world processes, including the processes modeled by GRNs [70, 93], it would be beneficial to characterize identifiability conditions under more general models, in order to increase the practicality of our theory and methods. We note that characterizing identifiability will be particularly challenging for non-additive noise models, where the noise depends on the process $X_{t}$. For example, the same marginal observations can be explained by a nonlinear additive noise SDE and a linear multiplicative noise SDE [18]. Our causal discovery results in Section 4.2 are also limited to additive noise. For this reason, it may be easier to consider identifiability for parameteric families of nonlinear drift under the additive noise model. Similarly, we note that it should be straightforward to extend experimental evaluation of APPEX to SDEs with nonlinear drift. Indeed, although the MLEs may not have a closed form solution, optimizers like BFGS [47, 20, 87] and neural networks [73] have been successfully used for SDE parameter estimation on real and synthetic data.
Third, our theory considers the infinite data setting, given the process’ true marginals $\{p_{i}\}_{i=0}^{N-1}$, and we have not studied asymptotic convergence rates. In particular, our identifiability result was established given continuously observed marginals, which corresponds to obtaining infinite samples from infinite marginals. Similarly, APPEX models the case where we obtain infinite samples from a finite set of marginals. We leave the asymptotics given the number of marginals $N$ and the number of observations per marginal $M$ for future work. Some theoretical considerations include the sample complexity of entropic optimal transport for trajectory inference, which requires asymptotically more samples per marginal as we decrease the entropic regularization (decreasing time granularity in the SDE setting) [29][Thm. 3], and the fact that the MLE of the AR(1) process from multiple trajectories is slightly biased [41]. We note that APPEX could be combined with previous methods that relax the marginal constraints (see [47]). This could improve estimates in low data settings, particularly in the biological setting where taking measurements destroys samples.
Finally, even in the asymptotic setting, the nonconvex nature of APPEX’s alternating projection optimization problem makes the prospect of proving convergence difficult. Although our experiments indicate that APPEX’s estimates approach the true solution, regardless of the initial guesses for the drift $A$ and observational diffusion $H$, we have not yet proved convergence due to the non-convexity of the optimization space. A formal proof of convergence and consistency would further strengthen its theoretical foundation.
Code Availability
The Python code for reproducing the experimental results and figures is available at https://github.com/guanton/APPEX. Analogous code in R is available at https://github.com/HydroML/X0isAllYouNeed.
Acknowledgements
The authors would like to thank United Therapeutics for supporting this research. GS and AW also acknowledge the support of the Burroughs Wellcome Fund.